OpenStudy (anonymous):

Evaluate. 1 – 3 • 7 + 4 • 2 A. –20 B. –28 C. –32 D. –12

OpenStudy (anonymous):

WAIT I THINK I KNOW THIS ONE

OpenStudy (anonymous):

u do to k12

OpenStudy (anonymous):

yah

OpenStudy (yolo_king):

remember first add then subtract and then multiply

OpenStudy (yolo_king):

What do you think is the answer?

OpenStudy (anonymous):

lost it sorry i suck ( *_* )

OpenStudy (anonymous):

b?

OpenStudy (yolo_king):

Nope, but it is a negative.

OpenStudy (yolo_king):

Wait lol there all negative

OpenStudy (yolo_king):

Not B.

OpenStudy (anonymous):

it really should say -28 but it didnt paste on correctly

OpenStudy (yolo_king):

7+4 = 11

OpenStudy (yolo_king):

1-11=-10

OpenStudy (anonymous):

so -10?

OpenStudy (yolo_king):

waity a minute .-. im doing it wrong

OpenStudy (yolo_king):

Ok the answer is -12

OpenStudy (anonymous):

dont you have to go by PEMDAS

OpenStudy (yolo_king):

mhmmm...

OpenStudy (anonymous):

and i got -28

OpenStudy (yolo_king):

.-.

OpenStudy (anonymous):

ohh okay thanks!

OpenStudy (yolo_king):

hmmmmm...

OpenStudy (yolo_king):

hold on

OpenStudy (yolo_king):

Lol I got it wrong Parentheses, Exponents, Multiplication, Division, Addition, Subtraction

OpenStudy (yolo_king):

I was mistaken

OpenStudy (anonymous):

thats fine

OpenStudy (anonymous):

so witch is it?

OpenStudy (yolo_king):

Oh your right 1 – 3 • 7 + 4 • 2 3*7 = 21 and 4*2 = 8 So we got 1-21+8 21+8=29 So 1-29 = -28 >.<

OpenStudy (yolo_king):

You are right :P

OpenStudy (anonymous):

thanks lol

OpenStudy (yolo_king):

:P thank you

OpenStudy (anonymous):

can you help with another?

OpenStudy (yolo_king):

Sure

OpenStudy (yolo_king):

Close the question and make a new post :) Don't forget to medal :D